9ył-49=0 i need three answers just help with how to do this one please
Answer by richwmiller(17219)

There is one real solution and you guessed it two complex solutions. Complex solutions always come in pairs.
y = (7/3)^(2/3)
y = (-7/3)^(2/3)=-0.879605+1.52352i
y = -(-1)^(1/3) (7/3)^(2/3)=-0.879605-1.52352i